# Self-Attention Decomposition for Training Free Diffusion Editing

### v1

@@ Abstract
Diffusion models achieve remarkable fidelity in image synthesis, yet precise control over their outputs for targeted editing remains challenging. A key step toward controllability is to identify interpretable directions in the model’s latent representations that correspond to semantic attributes. Existing approaches for finding interpretable directions typically rely on sampling large sets of images or training auxiliary networks, which limits efficiency. We propose an analytical method that derives semantic editing directions directly from the pretrained parameters of diffusion models, requiring neither additional data nor fine-tuning. Our insight is that self-attention weight matrices encode rich structural information about the data distribution learned during training. By computing the eigenvectors of these weight matrices, we obtain robust and interpretable editing directions. Experiments demonstrate that our method produces high-quality edits across multiple datasets while reducing editing time significantly by 60% over current benchmarks.
Fig. 1: Illustration of editing capabilities with our method. (a) The first row demonstrates the linear editing property as the perturbation strength $\alpha$ is varied. (b) The second row highlights the ability of our method to perform precise edits to facial attributes.
Index Terms—  Diffusion Model, Interpretibility Analysis
@@ 1 Introduction
Diffusion models [1] have emerged as the state-of-the-art for generative modeling, producing diverse, high-fidelity images by iteratively denoising random noise. A crucial application is to control their internal representations to support inference-time tasks such as semantic image editing. However, unlike single-pass generators, the multi-step, iterative nature of the diffusion process results in a unstructured latent space, where semantics are difficult to disentangle. To address this, prior works have explored various techniques for identifying interpretable editing directions. Some methods focus on semantic information within bottleneck features (e.g., the h-space framework) [2], while others operate directly on the noise latent space [3, 4]. Some other employ techniques like Riemannian geometry [5] and contrastive learning [6] to find robust editing vectors. However, these approaches face two key limitations. First, they typically rely on sampling a large number of images from the latent space, causing the discovered directions to be inherently tied to the sampled data distribution. This dependency can introduce a bias and limit the generalization to unseen images. Second, discovering such directions often involves computationally expensive operations [3, 5] or additional model training [6, 2]. These limitations substantially increase the editing time.
@@
    To the best of authors knowledge, we are the first to derive interpretable editing directions directly from the pretrained weights of diffusion models. We achieve this through an eigen-analysis of the self-attention weights, extracting eigenvectors without any additional training.
@@ 2.1 Preliminaries
$n^{*}=\arg\max_{n\in\mathbb{R}^{N\times d},|n_{i}\|_{2}=1}\|\Delta\mathrm{Attn}\|_{2}^{2}.$ (5)
Latent Whitening: Our derivation is based on the observation that the U-Net’s internal representations are whitened, i.e., their covariances are approximately identity during early denoising steps. This extends the empirically observed latent whitening effect, where intermediate feature covariances approach identity [11, 12] and enables the key analytical simplification $\mathbb{E}[X^{\top}X]\approx I$ for our intermediate matrices.
@@ 2.2 Weight space Eigen decomposition
To calculate the change in score matrix $\Delta L$ induced by the pertubration, while neglecting the $\mathcal{O}(\alpha^{2})$ term:
@@
Objective Function. Maximize $\Delta\mathrm{Attn}$ as defined in (5).
@@
$\bigl\|\Delta S\,V+S\,\Delta V\bigr\|_{F}^{2}$ $=\bigl\|\Delta S\,V\bigr\|_{F}^{2}+\bigl\|S\,\Delta V\bigr\|_{F}^{2}$ (12)
$\quad+2\,\mathrm{tr}\!\bigl((\Delta S\,V)^{\top}(S\,\Delta V)\bigr)$
Using whitening approximation (Sec. 2.1), which makes query–key and value perturbations nearly uncorrelated, we drop the cross term:
@@
From equations (8) and (16), we have $\Delta S$:
@@
With approximation (2.1), the latent-dependent covariance reduces to identity ($\mathbb{E}[VV^{\top}]\approx I$), isolating weight matrices:
@@
Since $\mathbb{E}[S^{\top}S]\approx I$, (Sec 2.1). The term then reduces to :
@@
$\mathbb{E}\bigl\|\Delta\mathrm{Attn}\bigr\|_{F}^{2}\approx\alpha^{2}n^{\top}\Bigl(W_{Q}^{\top}W_{Q}+W_{K}^{\top}W_{K}+W_{V}W_{V}^{\top}\Bigr)n$ (19)
@@
Now maximizing $n^{\top}C\,n$ subject to $\|n\|_{2}=1$ is a standard Rayleigh quotient problem. The optimal direction $n^{*}$ is the principal eigenvector of $C$. Subsequent eigenvectors provide orthogonal editing directions.
As shown in prior work[13, 3], editing vectors are most effective when applied at early denoising timesteps, before image content forms, where the whitening assumption $\mathbb{E}[ZZ^{\top}]\approx I$ holds. Following a similar strategy to prior works, we empirically choose the range $0.8T<t<0.5T$.
@@ 3.1 Qualitative Assessment
Figure 2 shows qualitative comparisons with benchmarks, demonstrating that our approach can edit diverse facial attributes while faithfully preserving the unedited regions compared to the baselines.
@@
Method Name Training-Editing Time Sample Independent Directions Localized Edit? Supervision Required? Images for Learning
@@ 3.2 Quantitative Analysis
In our quantitative analysis, we measure both image‐level fidelity and semantic alignment to the intended edits. We measure pixel- and structure-level consistency using SSIM and PSNR, identity preservation via FaceNet [20] cosine similarity, and semantic alignment with the Directional CLIP score [21]. For evaluation, we test 7 randomly selected editing directions on 15 image samples, resulting in 105 test cases. Table 2 compares our method and four baselines methods on the CelebA-HQ test set. Comparative results against existing benchmarks for other key evaluation metrics are reported in Table 1. Experiments show that our method reduces latency by 60% while consistently demonstrating superior results across baselines in both semantic alignment and image-level metrics.
@@ 4 Conclusion
In this paper, we introduce a training-free method for identifying editing directions in diffusion models via eigenanalysis of the self-attention weights in pretrained diffusion models. Our theoretical derivation in Sec. 2 demonstrates that the principal eigenvectors of the combined query-key-value weight matrices correspond to semantically meaningful editing directions, providing a closed-form solution to the image editing problem in diffusions. Experiments across diverse datasets show that these eigenvectors enable accurate and disentangled edits without any sampling or auxiliary trainings. Future work includes extending this weight-space analysis to multi-modal diffusion models, where cross-attention layers may reveal joint editing directions.

### v2

@@ Abstract
Diffusion models achieve remarkable fidelity in image synthesis, yet precise control over their outputs for targeted editing remains challenging. A key step toward controllability is to identify interpretable directions in the model’s latent representations that correspond to semantic attributes. Existing approaches for finding interpretable directions typically rely on sampling large sets of images or training auxiliary networks, which limits efficiency. We propose an analytical method that derives semantic editing directions directly from the pretrained parameters of diffusion models, requiring neither additional data nor fine-tuning. Our insight is that self-attention weight matrices encode rich structural information about the data distribution learned during training. By computing the eigenvectors of these weight matrices, we obtain robust and interpretable editing directions. Experiments demonstrate that our method produces high-quality edits across multiple datasets while reducing editing time by 60% over current benchmarks.
Fig. 1: Illustration of editing capabilities with our method. The first row demonstrates the linear editing property as the perturbation strength $\alpha$ is varied. The second row highlights the ability of our method to perform precise edits to facial attributes.
Index Terms—  Diffusion Model, Interpretability Analysis
@@ 1 Introduction
Diffusion models [6] have emerged as the state-of-the-art for generative modeling, producing diverse, high-fidelity images by iteratively denoising random noise. A crucial application is to control their internal representations to support inference-time tasks such as semantic image editing. However, unlike single-pass generators, the multi-step, iterative nature of the diffusion process results in a unstructured latent space, where semantics are difficult to disentangle. To address this, prior works have explored various techniques for identifying interpretable editing directions. Some methods focus on semantic information within bottleneck features (e.g., the h-space framework) [10], while others operate directly on the noise latent space [3, 20]. Some others employ techniques like Riemannian geometry [13] and contrastive learning [4] to find robust editing vectors. However, these approaches face two key limitations. First, they typically rely on sampling a large number of images from the latent space, causing the discovered directions to be inherently tied to the sampled data distribution. This dependency can introduce a bias and limit the generalization to unseen images. Second, discovering such directions often involves computationally expensive operations [3, 13] or additional model training [4, 10]. These limitations substantially increase the editing time.
@@
    To the best of the author’s knowledge, we are the first to derive interpretable editing directions directly from the pretrained weights of diffusion models. We achieve this through an eigen-analysis of the self-attention weights, extracting eigenvectors without any additional training.
@@ 2 Proposed Method
$n^{*}=\arg\max_{n\in\mathbb{R}^{N\times d},\|n_{i}\|_{2}=1}\|\Delta\mathrm{Attn}\|_{2}^{2}.$ (5)
Latent Whitening: Our derivation is based on the observation that U-Net intermediate representations during early diffusion denoising steps are approximately whitened, i.e., their feature covariances are close to identity. This extends prior empirical observations of latent whitening, where diffusion latents exhibit near-isotropic statistics due to injected Gaussian noise, particularly at high-noise timesteps [5, 1]. In this regime, which we empirically target ($0.8T\geq t\geq 0.5T$), the covariance structure of intermediate features becomes approximately diagonal, enabling the analytical simplification
$\mathbb{E}\!\left[Z^{\top}Z\right]\approx I.$ (6)
This isolates self-attention effects from the data distribution, enabling a closed-form eigen decomposition with semantically coherent directions across datasets and architectures.
@@ 2.1 Weight Space Eigen Decomposition
To calculate the change in score matrix $\Delta L$ induced by the perturbation, while neglecting the $\mathcal{O}(\alpha^{2})$ term:
@@
Objective Function: Maximize $\Delta\mathrm{Attn}$ as defined in Eq. (5).
@@
$\bigl\|\Delta S\,V+S\,\Delta V\bigr\|_{F}^{2}$ $=\bigl\|\Delta S\,V\bigr\|_{F}^{2}+\bigl\|S\,\Delta V\bigr\|_{F}^{2}$
$\quad+2\,\mathrm{tr}\!\bigl((\Delta S\,V)^{\top}(S\,\Delta V)\bigr).$ (13)
Under the whitening approximation(Sec. 2), perturbations to S and V become weakly correlated, allowing us to neglect the cross term in expectation.
@@
From Eqs. (9) and (10), we have $\Delta S$:
@@
With approximation (2), the latent-dependent covariance reduces to identity ($\mathbb{E}[VV^{\top}]\approx I$), isolating weight matrices:
@@
Since $\mathbb{E}[S^{\top}S]\approx I$, (Sec. 2). The term then reduces to
@@
$\mathbb{E}\bigl\|\Delta\mathrm{Attn}\bigr\|_{F}^{2}\approx\alpha^{2}n^{\top}\Bigl(W_{Q}^{\top}W_{Q}+W_{K}^{\top}W_{K}+W_{V}W_{V}^{\top}\Bigr)n.$ (20)
@@
Now maximizing $n^{\top}C\,n$ subject to $\|n\|_{2}=1$ is a standard Rayleigh quotient problem. The optimal direction $n^{*}$ is the principal eigenvector of $C$. Subsequent eigenvectors provide orthogonal editing directions. As shown in prior work [9, 3], editing vectors are most effective when applied at early denoising timesteps, before image content forms, where the whitening assumption $\mathbb{E}[ZZ^{\top}]\approx I$ holds. Following a similar strategy to prior works, we empirically choose the range $0.5T<t<0.8T$.
@@ 3.1 Qualitative Assessment
Method Name Editing Time Sample Independent Directions Localized Edit? Supervision Required? Images for Learning
@@
Figure 2 shows qualitative comparisons with benchmarks, demonstrating that our method edits diverse facial attributes while better preserving unedited regions than the baselines.
@@ 3.2 Quantitative Analysis
In our quantitative analysis, we measure both image‐level fidelity and semantic alignment to the intended edits. We measure pixel- and structure-level consistency using SSIM and PSNR, identity preservation via FaceNet [14] cosine similarity, and semantic alignment with the Directional CLIP score [8]. For evaluation, we test 15 semantic editing directions on 100 image samples, resulting in 1500 test cases. Table 2 compares our method and four baselines methods on the CelebA-HQ test set. Comparative results against existing benchmarks for other key evaluation metrics are reported in Table 1. Experiments show that our method reduces latency by 60% while consistently demonstrating superior results across baselines in both semantic alignment and image-level metrics.
@@ 4 Conclusion
In this paper, we introduce a training-free method for identifying editing directions in diffusion models via eigenanalysis of the self-attention weights in pretrained diffusion models. Our theoretical derivation in Sec. 2 demonstrates that the principal eigenvectors of the combined query-key-value weight matrices correspond to semantically meaningful editing directions, providing a closed-form solution to the image editing problem in diffusions. Future work includes extending this weight-space analysis to multi-modal diffusion models, where cross-attention layers may reveal joint editing directions. Extending our eigen-analysis to the cross-attention layers of recent DiTs for joint text–visual control is a direction for future work, particularly to study robustness in higher-dimensional latent spaces.